OpenStudy (anonymous):

Among all pairs of numbers whose difference is 10, find a pair whose product is as small as possible. What is the minimum product? a. the pairs of numbers whose difference is 10 and whose product is as small as possible is___ b. The minimum product is___

OpenStudy (anonymous):

bet it is -5 and 5

OpenStudy (anonymous):

call one number x, the other x + 10 then their product is \[x(x+10)=x^2+10x\]

OpenStudy (anonymous):

this thing has a minimum value at \[x=-\frac{b}{2a}=-\frac{10}{2}=-5\]

OpenStudy (anonymous):

x = -5, x+10 = 5 and their product is -25